The Kaufman Assessment Batteryfor Children is designed to measure achievement and intelligence with a special emphasis on nonverbal intelligence Its global measures, such as its Sequential Processing score, are scaled to have a mean of 100 and a standard deviation of 15. Assume that the Sequential Processing score has a normal distribution. (a) Find a value such that divides the children with the highest 10% of the scores from those with the lower 90%Can find Z from the table at 10% (0.10) (ie. 0.1003) this is2 Z-Y-u need to calculate Y Table A.1 (Continued) 0.33 0.3707 074 0.2290 1.15 0.1251 15e 0.04 03 0.3009 0.75 02200 1.16 0.1230 1570.052 0.3 0.3032 0.70 0.223 1.17 0.1210 158 0067 0.30 03594 0.77 0220 1.18 0.1100 1.59 0.059 037 03557 078 02177 1.19 0.1170 1eo。сее 038 0.3520 079 0.21481.20 0.1151 101 0.0637 030 0.3483 080 02119 121 0.1131 102 00ae 0A0 03446 081 02090 122 ·1112 ocele 041 0.3400 082 02001 1.23 .1093 104 05 042 0.3372 0 83 020S 1.24 01075 160 00495 430.3330 084 02005 125 1050 1.00 0048s 440.300 085 0.1977 12 1038 17 0047 045 02064 ose, o.1949 1.27。1000 ,66 00406 046 0.320.8 01922 12 0 1000 1.00 00435 04 0.31 088 0.1804 129 009es 1.70 044 0 48 0 3150 0.a o.1507 1.30 171 0043 40 03121 0.90 0.1841 131 0.01 1.72 00427 50 0.3080 091 0.1814 132 004 173 004 00 300 092 0.1768 13 0 0018 1.74 0040 o 52 o 3015 0930.1762 1.34 0.001 1.75 0040 053 0 24 094 01730 1.3s a os1.70 00302 54 02040 095 01711 13 0ee 1.7 0084 ose. 02912。96 0.1685 1.37 00853 1,78 00075
Answer #1

a)from table for 90 th percentile ; critical z =1.28

therefore corresponding value =mean+z*std deviation =100+1.28*15=119.20